HELP PLEASE ITS ON ATOMIC MASS AND NUCLEAR PROCESSES

Astronomy homework - due Friday, Express all yourthanks

˝
Answer #1

1) 13N7

0p1

4He2


2)a) 235U92

b) Mass disappeared = 2(1.007276 + 1.008665 + 0.000549)u = 4.03298u


3) Difference in mass = 6(1.007276 + 1.008665 + 0.000549)u - 12u = 12.09894u - 12u = 0.09894u

So, Mass of the produced carbon atom is less.


4) 14N7 + 1n0 = 14C6 + 1p1